### v3

# Probing For Non-Gravitational Long-Range Dark Matter Interactions

Abstract
Dark matter remains a mystery in fundamental physics. The only evidence for dark matter’s existence is from gravitational interactions. We constructed a precision torsion balance experiment to search for non-gravitational, long-range interactions between ordinary matter in our lab and the Milky Way’s dark matter. We find no evidence of such interaction and set strict upper bounds on its strength. These results suggest that dark matter only interacts gravitationally over long distances and constrains a variety of dark matter theories.
Introduction - Dark matter has remained a persistent mystery in fundamental physics. Decades of cosmological observations – including galactic rotation curves Salucci (2019), large-scale structure surveys Collaboration (2024, 2025), and measurements of the cosmic microwave background Collaboration (2020) – have established that dark matter constitutes approximately 26% of the universe’s mass-energy content, far greater than the  5% contribution from ordinary matter. Collaboration (2020) The luminous matter of the Milky Way galaxy is embedded in a halo of dark matter that makes up most of the galaxy’s mass. Hunt and Vasiliev (2025) Despite this apparent abundance, the true identity of dark matter remains elusive and it is still unknown whether dark matter interacts in another way beyond gravity.
Figure 1: A schematic of the experimental setup in the galactic frame. The experiment searches for differential accelerations of two different materials towards the dark matter in the Milky Way galaxy. Adapted from ESA/Gaia/DPAC, Stefan Payne-Wardenaar. ESA/Gaia/DPAC (2025)
The prevailing paradigm envisions dark matter as a new class of non-baryonic, weakly-interacting particles Bertone (2010); Arcadi et al. (2018). However, every effort at detecting and producing particle dark matter has come up empty. Bozorgnia et al. (2024) This invites serious consideration of nontraditional possibilities including: primordial black holes Villanueva-Domingo et al. (2021), sterile neutrinos Boyarsky et al. (2019), and entirely undiscovered phenomena.
Figure 2: Schematic of the rotating torsion balance apparatus with embedded picture of the pendulum held before installation. The thermal shields, vacuum chamber, and magnetic shields have been cut away to show the torsion balance held inside. The torsion pendulum has four test bodies made of aluminum and four of beryllium in a dipole orientation. The pendulum is suspended by a quartz torsion fiber attached to a magnetic swing damper whose position can be actuated by the positioning stage. The vacuum chamber, torsion balance, autocollimator, and magnetic shield are suspended from the air-bearing turntable. The rotating tilt sensors are used to level the turntable by actuating the thermal expansion feet. The angle of the pendulum relative to the vacuum chamber is measured by the autocollimator.
We constructed a laboratory experiment to search for non-gravitational, long-range interactions between ordinary matter and the Milky Way’s dark matter. We measured the acceleration difference between two different materials directed towards the center of the Milky Way galaxy, as shown in Figure 1. Resolving such minuscule differential accelerations required development of an ultra-sensitive torsion balance apparatus.
Apparatus - Torsion balances are highly sensitive mechanical systems that have long been used for fundamental physics experiments. Adelberger et al. (2009) They are formed by a pendulum suspended from a thin fiber that acts as a weak torsional spring.
Our torsion pendulum consisted of a light-weight frame that held four test bodies made of beryllium and four made of aluminum. This formed a composition dipole oriented perpendicular to the torsion fiber (Figure 2). The pendulum was designed with a high degree of symmetry to minimize coupling to gravitational gradients and other known interactions, such as electrostatics. The external surfaces were coated with gold to reduce surface charges. The pendulum was suspended from a 22-µm thick, 1-m long fused quartz fiber Shaw et al. (2022), which gave the balance a torsional period of 1472 seconds and a ring down time of $\sim$2 years.
The torsion balance was held inside a vacuum chamber maintained at $0.2~\text{mPa}$ by an ion pump. At low frequencies ($1~\text{mHz}$), internal thermal noise Saulson (1990) of the torsion fiber limited the performance of the balance. The angle of the pendulum relative to the vacuum chamber was measured with an autocollimator. The autocollimator reflected a laser beam off one of the four mirrors attached to the pendulum. Table 1 displays some of the relevant parameters of the torsion balance apparatus.
To avoid slowly varying noise (i.e. 1/f-noise), we modulated the signal of interest by suspending the entire apparatus from a high-precision air-bearing turntable that was set to continuously rotate at 2.85 mrad/s (2/3 of the torsional resonance frequency). The turntable was located on the upper floor of our laboratory with the rest of the apparatus extending to the floor below. The turntable’s angle encoder was locked to an external clock which maintained a constant rotation rate by varying the torque applied with an eddy current motor. Shaw (2023)
Misalignment of the rotation axis of the turntable and the torsion fiber axis (i.e. local vertical) can cause a parasitic once per turntable-revolution modulation by varying the torsion fiber attachment, translating the autocollimator beam, and/or changing the electrostatic environment around the pendulum. Such a modulation would not be correlated with the dark matter signal but could be a significant noise source. To mitigate this, we devised a control system that processed the signal from a pair of co-rotating tilt sensors to control the temperature of two of the three feet on which the turntable rests.
Figure 3: Once per turntable revolution differential acceleration amplitudes along with the corresponding fits to the galactic basis functions. The mean of each data set has been removed. Each two-day long segment was independently fit. Short gaps in the data were due to local construction activity. Data for sidereal days 189 to 285 was taken in 2024, while the remaining data was taken in 2025. The right panel shows the histogram of the differential acceleration amplitudes, the bottom panel shows the orientations of the pendulum with respect to the vacuum chamber over the data run. The zoom shows a two-day long segment and it’s corresponding fit to the galactic basis functions.
Thermal expansion of the feet aligned the turntable rotation axis to local vertical. Shaw (2023)
The pendulum was built from strictly non-magnetic materials and machined with ceramic tools to avoid ferromagnetic impurities. Additionally, the pendulum was surrounded by three layers of co-rotating magnetic shielding and one non-rotating shield to minimize the effects of external magnetic fields. The inner-most magnetic shield was carefully demagnetized to avoid eddy current damping on the pendulum which preserved the high quality factor and low thermal noise.
To reduce effects of temperature variations, the vacuum chamber was surrounded by a co-rotating sheet metal shield, a concentric lab-fixed shield made of aluminum and insulating foam, and an enclosure made of construction insulation. The torsion fiber was also surrounded by a thermally-isolated thick-walled copper tube. Finally, the experiment was located in an underground laboratory which was temperature controlled to $\pm 25~\text{mK}$ daily variation with a non-cycling air conditioner.
Apparatus
Parameter Value
Spring Constant $\kappa$ $7.1\times 10^{-10}$ N m/rad
Quality Factor $Q$ $1.1\times 10^{5}$
Mass of test body set $m$ $19.4$ g
Moment of inertia II $3.78\times 10^{-5}$ kg $\text{m}^{2}$
Resonant frequency $\omega_{0}$ $2\pi\times 0.69$ mHz
Turntable frequency $\omega_{TT}$ $2\pi\times 0.46$ mHz
Lever-arm $r$ $1.9$ cm
Fiber Length $l$ $0.94$ m
Table 1: Relevant parameters of the torsion balance. The lever-arm $r$ is the radial distance from the pendulum’s fiber axis to the center of mass of each set of test bodies.
Analysis - The experiment recorded data from July 7, 2024 to December 25, 2025. We rotated the pendulum $180^{\circ}$ with respect to the vacuum chamber on July 29, 2024, February 15, 2025, May 1, 2025, and September 29, 2025 to reduce systematic effects that may be associated with the pendulum orientation relative to the vacuum chamber (such as tilt and electrostatics). Rotating the pendulum changes the sign of the science signal while maintaining most systematics; these pendulum reorientations were accounted for in the analysis.
The duty cycle of the apparatus was severely restricted due to nearby construction activity and an unexpected hardware failure. Although the data spans 537 days, only 244 days (45%) were free of external disturbances and included in the data set. The measured torsion balance angle was calibrated with a programmed turntable speed change (see Supplementary Material). The angle was converted to differential acceleration by inverting the harmonic oscillator response of the pendulum with a time-domain filter and scaling by the relevant fixed parameters. An additional low-pass filter (1 mHz third-order Butterworth) was applied to minimize the effect of high frequency noise. We corrected for the response of this filter at the turntable frequency (0.4% attenuation, $\sim 54^{\circ}$ phase). The differential acceleration time series was split into two turntable-rotation long sets. Each set was fit to a linear combination of sinusoidal functions with frequencies of harmonics of the turntable angle ($\omega_{TT}$, $2\omega_{TT}$, $3\omega_{TT}$, etc.) and harmonics of the torsional resonance ($\omega_{0}$, $2\omega_{0}$, $3\omega_{0}$, etc.). The full fit function is shown in the Supplementary Material.
Figure 4: Measured accelerations from each two-day long segment towards the galactic center (in-phase) and the orthogonal direction (out-of-phase). The blue indicates the points taken at a pendulum orientation of $0^{\circ}$ and yellow indicates $180^{\circ}$. The red cross and circle, respectively, indicate the mean and corresponding uncertainty. The top and right panels show the histograms for the in-phase and out-of-phase accelerations, respectively.
The instrument experienced occasional transients due to earthquakes, nearby construction activity and other environmental disturbances. We discarded data sets that fell under two criteria: misfit-squared exceeding the 99th percentile of the corresponding $\chi^{2}$-distribution, and a once per turntable-revolution ($\omega_{TT}$}) acceleration component outside of the 95th percentile of the component’s distribution. We subtracted the mean $\omega_{TT}$}-acceleration amplitude of each data set leaving only the time varying components. Figure 3 shows the surviving once per turntable revolution acceleration amplitudes. The $\omega_{TT}$-acceleration was then extracted in two-day long segments. Segments that had less than a quarter of a sidereal day of data were discarded. Each segment was fit to galactic basis functions. These functions consisted of the location of the center of the Milky Way (RA: $18^{\text{h}}$ Dec: $29^{\circ}$) projected into the horizontal plane of the lab and an orthogonally oriented function. These functions modulate once per sidereal day due to the rotation of the Earth and the location of our lab at $\sim 48^{\circ}$ latitude. The location of the galactic center and the necessary coordinate transformations were computed using the Astropy Astropy Collaboration and Astropy Project Contributors (2022) libraries. This fitting yielded 56 points for the $0^{\circ}$ pendulum orientation and 66 points for $180^{\circ}$ for a total of 122 independent measurements.
Results - The series of measured differential accelerations toward the galactic center is shown in Figure 4. Measuring the out-of-phase component verifies the sensitivity of the experiment as the signal of interest would only appear as a non-zero in-phase component. We take the mean and corresponding uncertainty calculated from the scatter of the measurements to give a final measure of the differential acceleration:
$\Delta a_{DM,Be-Al}=0.5\pm 0.6\text{ fm/s}^{2}\quad(1\sigma)$ (1)
A positive $\Delta a$ would correspond to beryllium attracting more towards dark matter. The average acceleration for the 0◦ and 180◦ orientations are individually consistent with zero.
The gravitational acceleration on Earth towards the galactic dark matter is $\geq$50 $\text{pm/s}^{2}$. Stubbs (1993) If the galactic dark matter causes an acceleration that distinguishes between beryllium and aluminum, the interaction must be smaller than the gravitational acceleration by a factor of:
$\eta_{DM,Be-Al}\leq 2.4\times 10^{-5}\quad(95\%\text{-confidence}).$ (2)
These constraints are a factor of four improvement over previous measurements. Wagner et al. (2012)
Discussion - This observation can be applied to a variety of theoretical scenarios; many of which depend on the use of different test body materials. We chose to use aluminum and beryllium test bodies to maximize sensitivity to the leading theories (discussed below). Yet, our observations limit any long-range interaction between dark matter and normal matter that does not couple identically to aluminum and beryllium. Stubbs (1993) Other experiments have made different material choices and thus restrict these finely-tuned scenarios as well. Wagner et al. (2012)
There is no underlying symmetries in the Standard Model that require conservation of baryon number (B) and lepton number (L). Broussard et al. (2025); Dev et al. (2024); Perez et al. (2022); Cai et al. (2018) Yet, global violation of these conservations have not been observed. This has led to speculation that there may exist a B- or L-dependent fifth force. Such fifth forces have been stringently tested for normal matter. Touboul et al. (2022) Our results extend these tests to include dark matter-normal matter interactions.
Grand Unified Theories De Boer (1994) attempt to unify the electromagnetic, weak, and strong forces of nature into a single force. In many of these theories, B and L numbers are not conserved; instead the difference (B-L) is the conserved quantity. Buchmüller et al. (1991); Basso et al. (2009) This conservation has many possible consequences for dark matter including dark matter having B-L charge. Rodejohann and Yaguna (2015); Guo et al. (2015) Our experiment is uniquely designed to search for interactions to this B-L charged dark matter; the test body materials were chosen to maximize their difference in B-L charge. There is ambiguity in what B-L charge dark matter could have. Regardless, our results restrict the strength of such an interaction.
There are a lacking number of dedicated tests of quantum gravity theories. Despite this, we can test a few aspects of quantum gravity in laboratory experiments including the Weak Gravity Conjecture. Harlow et al. (2023); Arkani-Hamed et al. (2007) A statement of the conjecture is that gravity has to be the weakest force of nature to have a consistent theory of quantum gravity. Our observations show that the conjecture is still valid for interactions between dark matter and normal matter.
The Equivalence Principle is a foundational aspect of our geometric understanding of gravity (i.e. General Relativity). The principle states that gravity is independent of the composition of objects. This have been tested extensively for a variety of materials and distance scales. Touboul et al. (2022); Wagner et al. (2012) However, these tests use normal matter as the source mass making them blind to any violation associated with dark matter. Our results directly test the Equivalence Principle towards dark matter at galactic scales and set limits on the Eötvös parameter for beryllium and aluminum towards dark matter ($\eta_{DM,Be-Al}$ in Equation 2).
Conclusion - We constructed a rotating torsion balance to search for a long-range interaction between normal matter and dark matter. We observe no evidence for such an interaction and set strict limits on its possible strength. Although this constrains a variety of theoretical possibilities for dark matter, our observation is model independent. Any proposed dark matter interaction that does not couple identically to beryllium and aluminum is constrained by this observation.
There are very few laboratory experiments that directly probe dark matter without assuming a theoretical model. Our experiment achieved this on a galactic scale while being located on Earth. This observation furthers our understanding of not just dark matter but also the nature of gravity and fundamental symmetries of the universe.
Data Availability - Raw data will be shared upon request. The fit amplitudes and analysis code can be found at https://github.com/EotWash/EP-Analysis.
Acknowledgements - This work was supported by funding from the NSF under Awards PHY-1607385, PHY-1607391, PHY-1912380, and PHY-1912514.